OpenStudy (anonymous):

x^2=3y (x and y represent rect. coord.) write the equation using polar coordinates

OpenStudy (anonymous):

x=\[p \cos (\phi)*psin(\theta)\] y=\[psin (\phi)*psin(\theta)\]